### v2

# Ban&Pick: Ehancing Performance and Efficiency of MoE-LLMs via Smarter Routing

Abstract
Sparse Mixture-of-Experts (MoE) has become a key architecture for scaling large language models (LLMs) efficiently. Recent fine‑grained MoE designs introduce hundreds of experts per layer, with multiple experts activated per token, enabling stronger specialization. However, during pre‑training, routers are optimized mainly for stability and robustness: they converge prematurely and enforce balanced usage, limiting the full potential of model performance and efficiency at inference. In this work, we uncover two overlooked issues: (i) a few highly influential experts are underutilized due to premature and balanced routing decisions; and (ii) enforcing a fixed number of active experts per token introduces substantial redundancy. Instead of retraining models or redesigning MoE architectures, we introduce Ban&Pick, a post-training, plug-and-play strategy for smarter routing. Pick discovers and reinforces key experts—a small group with outsized impact on performance—leading to notable accuracy gains across domains. Ban further dynamically prunes redundant experts based on layer and token sensitivity, delivering faster inference with minimal accuracy loss. Experiments on fine-grained MoE-LLMs (DeepSeek, Qwen3) across math, code, and general reasoning benchmarks demonstrate that Ban&Pick delivers free performance gains and inference acceleration without retraining or architectural changes. For instance, on Qwen3-30B-A3B, it improves accuracy from 80.67 to 84.66 on AIME2024 and from 65.66 to 68.18 on GPQA-Diamond, while accelerating inference by 1.25× under the vLLM.
1 Introduction
Large language models (LLMs) have achieved remarkable progress (DBLP:conf/iclr/ZhouWLSLQLJSZ024, ), yet the ever-growing demand for higher capacity presents significant challenges for efficiency. As a promising solution, Mixture-of-Experts (MoE) architectures scale parameter count without proportionally increasing computation, by activating a small subset of experts for each input (artetxe-etal-2022-efficient, ).
Recent advances have pushed MoE towards a fine-grained design (dai-etal-2024-deepseekmoe, ), where each layer contains a large number of relatively small experts. For example, DeepSeek-v2.5 (deepseekai2024deepseekv2strongeconomicalefficient, ) employs 160 experts per layer, while Qwen3-MoE (yang2025qwen3technicalreport, ) adopts 128 experts per layer. Unlike coarse-grained MoE (jiang2024mixtral, ), which typically consists of a few large experts derived from the same base model, fine-grained MoE initializes each expert independently and randomly. This design promotes specialization: experts gradually develop distinct skills, excelling in areas such as math, code, or general reasoning. Such strong differentiation suggests that fine-grained MoE has the potential to leverage expert diversity more effectively.
However, we argue that a critical but overlooked phenomenon exists: the routing strategies learned during pre-training cannot fully exploit this specialization. Specifically, we empirically discover that current routing strategies inadvertently limit the potential to leverage the most impactful experts while simultaneously introducing redundancy where many selected experts contribute minimally to the final output. Instead of attempting to redesign training objectives or expert architectures, we take a post-training perspective, aiming to unlock this untapped potential efficiently and effectively.
Delving into this question, we surprisingly find that within the most frequently selected experts of fine‑grained MoE models, a small fraction plays an outsized role: simply forcing these experts to be chosen for every input can noticeably boost accuracy on certain datasets. Motivated by this striking observation, we conduct an empirical analysis of expert usage across domains and discover that each domain consistently relies on its own set of frequently selected experts, which we refer to as domain‑specialized experts. Yet, a closer examination further shows that only a tight subset of these experts has a decisive impact on the final logits distribution, while the influence of the others remains marginal. We term these most critical ones key experts. Our Pick module is designed to reinforce their influence during routing, getting performance gains by emphasizing these key experts.
In parallel, we aim to alleviate redundancy in expert activation. Under the constraint that a fixed number of experts must be selected for each token, many activated experts contribute little to the final output, reflecting the redundancy introduced by less relevant experts. Our Ban module jointly considers layer sensitivity and token sensitivity to assess the redundancy in expert selection, and applies dynamic pruning of experts during inference. In this way, Ban reduces unnecessary computation and achieves significant inference acceleration with only minor accuracy loss.
Attempting to bridge the gap between routing strategies learned during pre-training—which prioritizes stability and balance—and optimal inference-time expert utilization, Pick and Ban empower truly impactful experts to contribute more while reducing participation from low-contribution experts:
  • Pick amplifies excellence by identifying experts that demonstrate superior performance on specific domains and strategically increasing their influence during inference, maximizing effectiveness.
  • Ban eliminates inefficiency by dynamically excluding experts that contribute minimally during inference, reducing computational costs while preserving essential capabilities.
Together, we introduce Ban&Pick, a unified framework that optimizes expert utilization in inference to improve both the accuracy and efficiency of fine-grained MoE without retraining or architectural changes. Evaluated on five widely used datasets spanning math, code, and general reasoning, the results demonstrate that on the Qwen3 series models, applying Pick alone brings an average 2.83% performance improvement. By further incorporating Ban, the joint Ban&Pick strategy achieves an average 1.41% improvement while simultaneously accelerating inference by 1.25× under the vLLM.
2 Related Works
Enhancing Expert Utilization in MoE.
Most existing work attempts to improve expert utilization efficiency in MoE models by redesigning training objectives or expert architectures during pre-training (yan2025tcmoe, ; huang-etal-2024-harder, ; jin2024moeacceleratingmixtureofexpertsmethods, ). However, while these approaches are insightful, they often sacrifice training stability for flexibility. Recently, (wang2025expertsneedsteeringthinking, ) proposes a post-training method designed for MoE models that employ the <think> token, enhancing reasoning by amplifying the weights of Cognitive Experts.
Dynamic Expert Pruning of MoE.
Dynamic expert pruning has emerged as a key technique for accelerating MoE inference. (lu-etal-2024-experts, ) proposed a pruning strategy that simply skips experts with less weights. Building on this, (huang2024mc, ) introduced a critical token protection mechanism. However, the substantial sensitivity differences across layers of MoE models and across tokens in long chain‑of‑thought reasoning are often overlooked, leaving room for further optimization.
3 Empirical Findings on Routing Limitations in Fine-grained MoE
Figure 1: The intuition and empirical findings on expert utilization. Left (Pick): forcibly activating key experts (e.g., E43) improves accuracy on Winogrande (sakaguchi2021winogrande, ). Right (Ban): reducing experts from 6 to 5 cuts 16.7% computation without accuracy loss.
3.1 Insights from Pre-training Patterns
Recent evidence from OLMoE (muennighoff2025olmoeopenmixtureofexpertslanguage, ) and OpenMoE (xue2024openmoe, ) suggests that during pre-training, routers converge much earlier than the experts. In particular, OLMoE reports that after only 1% of pre-training, around 60% of the routing decisions have already saturated, while OpenMoE finds that by 40% of pre-training, about 80% of the routing has stabilized. Such early fixation implies that routing decisions are largely determined before the experts have sufficiently matured. Moreover, the widely used balancing loss (lepikhin2020gshardscalinggiantmodels, ), while effective at preventing collapse to just a few experts, may also discourage tokens from concentrating on the most impactful experts. This implies that the current pre-training paradigm for MoE makes a trade-off: in exchange for training stability, it locks routers into decisions before experts have fully matured. Combined with balancing constraints, this underutilizes high-potential experts and inflats the role of those that contribute little at inference. This presents an opportunity from a post-training perspective:
Can we recover this lost potential—amplifying the influence of the most impactful experts while trimming away redundancy—to both boost performance and accelerate inference?
3.2 Expert Utilization Inefficiencies and Optimization Potential
To validate the inefficiencies in expert untilization and better understand the optimization potential, we conducted two simple yet revealing experiments with the Deepseek‑V2‑Lite‑Chat model (activates 6 experts out of 64). Figure 1 demonstrates how we adjust expert utilization and its effects.
(1) Underutilization of the most impactful experts: we examined expert selection frequency on the C4 (dodge2021documentinglargewebtextcorpora, ) calibration set, identified the top-3 most frequently selected experts in each layer, and then conducted a grid search: each of these experts was manually added in addition to the original routing (E43 in Figure 1), and the resulting accuracy on Winogrande benchmark was recorded. The left part highlights a few experts contribute disproportionately: enhancing Layer‑1 Expert‑43 improved accuracy by +1.81%, L10E33 by +1.65%, and L25E11 by +1.26%. When these three were jointly enhanced, accuracy increased by +2.72% (71.67% $\rightarrow$ 74.39%), while most other high-frequency experts had negligible or noisy effects, indicating that certain impactful experts are insufficiently leveraged in the original routing strategy. We refer to these three experts as key experts.
(2) Computational waste from low-contribution experts: we directly reduced the number of selected experts per token from 6 to 5 and measured the resulting accuracy. Remarkably, as shown in the right side of Figure 1, this pruning reduced computation by 16.7% and surprisingly preserved full accuracy on Winogrande (71.67%), demonstrating redundancy in the original expert utilization.
Together, these experiments validate the inefficiencies in expert utilization predicted by pre-training analysis, and support the intuition of our proposed methods—Pick and Ban, respectively, which aim to optimize expert selection beyond pre-training constraints for better inference-time performance. This empirical study motivates us to design a unified framework that can pick key experts while banning redundant ones, aiming at improving both accuracy and efficiency.
4 Pick
In this section, we begin with an analysis of expert specialization in fine‑grained MoE models to verify the existence of domain-specialized experts, which forms the foundation for key experts to effectively function. Building on this analysis, we then introduce our Pick method in two steps: (i) identifying key experts from the set of domain‑specialized experts, and (ii) designing strategies to enhance these key experts so that their impact on model performance can be effectively amplified.
4.1 Expert Specialization in Fine-Grained MoE
Figure 2: Expert specialization in fine-grained MoE (Qwen3‑30B‑A3B). Left: expert usage frequency across three tasks (math, general, code). Right: token word clouds for 3 high-frequency experts.
First, we take a closer look at the degree of expert specialization in fine‑grained MoE models. Specifically, we analyze the Qwen3‑30B‑A3B model and record expert usage frequencies on task-related benchmarks: math tasks (GSM8K + Math-500), code tasks (MBPP (austin2021program, ) + Livecodebench), and a general task (GPQA-Diamond). The statistics cover all tokens in both the prefill and decode stages. For each type of task, we plot the per‑layer expert usage frequency and show 3 representative layers in Figure 2 (complete results in Appendix A).
Our analysis reveals that in each layer there exist experts whose usage frequency is significantly higher than average–often up to 6 times higher. Moreover, in most layers these high‑frequency experts vary across tasks. For instance, L16E95 is frequently activated only on math tasks, whereas L29E91 is almost exclusively activated on code tasks with an extremely high activation rate of 66.9%. In contrast, a small number of experts such as L4E48 remain highly active across all categories of tasks.
These task‑dependent high‑frequency experts suggest the existence of domain‑specialized experts. To validate this hypothesis, we further examine the functional roles of these experts by analyzing which tokens are most often routed to them. Taking the three aforementioned experts as examples, we visualize the tokens they frequently process using word clouds (after removing empty or semantically meaningless tokens for readability). The results show that the math‑specialized expert L16E95 is strongly associated with digits, operators, and math terminology; the code‑specialized expert L29E91 is associated with common programming tokens; and the general expert L4E48 is closely tied to branch tokens (zhong2025hybrimoehybridcpugpuscheduling, ), which play a critical role in chain‑of‑thought across diverse tasks.
These observations provide empirical evidence that expert specialization in fine‑grained MoE models is highly pronounced. Different tasks rely heavily on different domain‑specialized experts, while a small subset of general experts remain consistently important across domains.
4.2 Picking Key Experts from Domain‑Specialized Experts
As shown previously, almost every layer contains domain‑specialized experts that are activated with notably high frequency. The key issue, however, is whether all of these experts contribute equally to model’s performance, or only a small subset is truly essential. To investigate this, we evaluate the importance of each expert by measuring the shift in the logits distribution when it is removed.
Concretely, taking math tasks as an example, we use the domain‑specialized experts identified in the previous analysis as candidate experts in each layer. We then run the model on 1000 randomly sampled instances from the MathQA (amini-etal-2019-mathqa, ) dataset. For each sample, we record the output logits of the original model, and then repeat the inference while pruning one candidate expert at a time. The perturbation of the output distribution is quantified using the KL divergence between the probability distributions of the Top‑1000 tokens from the original and pruned models (see Appendix B for implementation details).
The results are summarized in Figure 4. Surprisingly, despite their high activation frequency, most domain‑specialized experts cause only negligible shifts in the output distribution when pruned. In contrast, a small number of experts stand out, with three of them inducing particularly large changes. Among these, expert L9E18 is the most striking, leading to a divergence of about $0.16$. These findings suggest that only a few domain‑specialized experts truly have a decisive influence on the model’s behavior, and we hypothesize that they are the key experts we aim to identify.
To test this hypothesis, we construct a failure set consisting of the Math‑500 (lightman2023lets, ) problems that the base model originally answers incorrectly (24 in total). For each candidate expert, we forcibly include it during routing and then re‑evaluate accuracy on this failure set. As shown in Figure 4 (right vertical axis), only those experts that induce substantial distribution shifts also yield meaningful accuracy gains when enhanced, whereas others contribute little to no improvement. In particular, expert L9E18, which caused the largest divergence, also increases the number of correct answers by 5 out of 24 when included. This provides strong evidence of the close correspondence between distribution shifts and functional importance, and offers a principled way to distinguish key experts from the many domain‑specialized experts.
4.3 Enhancing Key Experts
Figure 3: Impact of pruning (measured by KL divergence, left axis) and enhancing (by accuracy gain, right axis) for domain‑specialized experts.
Figure 4: Comparison of five designed enhancement methods for key experts, evaluated by accuracy on three widely used math benchmarks.
After identifying key experts, the remaining challenge is to determine how to leverage them most effectively. We take math tasks as a running example and design five enhancement strategies:
  1. A.
    Forced Addition: Whenever a key expert is not selected, it is added alongside the chosen experts.
  2. B.
    Forced Replacement: Similar to forced addition, but the key expert replaces the selected expert with the lowest routing weight, maintaining the total number of selected experts.
  3. C.
    Range-based Addition: If a key expert is not selected (not in the top-$k$ candidates) but appears within the router’s top-$2k$ candidates, it is forcibly added.
  4. D.
    Range-based Replacement: Under the same condition as range-based addition, the key expert replaces the selected expert with the lowest routing weight.
  5. E.
    Logit Bias: Before routing, a fixed bias is added to the routing logits of key experts. In our implementation, the bias is set to $20\%$ of the average weight among the top-$k$ selected experts.
We evaluate these methods on AIME2024, AIME2025, and Math‑500. For the two AIME datasets, results are averaged over five runs to mitigate small‑sample variance. Figure 4 summarizes the outcomes. Overall, strategy A brings consistent but modest gains. Strategy B, which forces replacement regardless of probability, is unstable and even reduces accuracy on AIME2024 and Math‑500, indicating that overly aggressive intervention can be harmful. By contrast, strategies C and D—enhancing only when the key expert is already close to the router’s top candidates—deliver the strongest improvements across all benchmarks. The difference between C and D is minor, but since D avoids expanding the number of active experts, it is preferable in practice and used in our experiments. Strategy E shows less stable performance, as the bias magnitude is sensitive and difficult to tune.
5 Ban
Our dynamic expert pruning method is motivated by two key observations: 1) Different layers in an MoE model exhibit highly diverse sensitivity to expert pruning and thus should be treated differently. 2) During reasoning, the token-wise routing distributions also vary significantly: when most of the routing weight is concentrated on a few top experts, expert pruning causes minimal loss; in contrast, when the weights are more evenly distributed, expert pruning becomes more harmful.
In the following, we first use Qwen3-30B-A3B as an example to illustrate the differences in layer sensitivity and token sensitivity, and then introduce our proposed dynamic pruning method.
Layer Sensitivity.
We measure layer-wise pruning sensitivity by reducing the number of selected experts from $k=8$ to $k=3$ in each layer, and computing the Top‑1000 KL divergence between the output logits of the original and pruned models on MathQA calibration set. The results (Figure 5a) reveal substantial differences across layers: while some layers, especially at the begin and end of the model, are highly sensitive to pruning, many middle layers are much more robust. This large variance highlights the necessity of treating different layers differently when designing pruning strategies.
Token Sensitivity.
We further analyze token‑level pruning sensitivity in long reasoning processes by examining how routing weights are distributed across experts. To quantify this, we compute the ratio between the cumulative routing weights of the top‑3 and top‑8 experts, averaged across all layers. Figure 5b plots these ratios for the first 200 generated tokens in a reasoning process from one GPQA‑Diamond sample. The results reveal substantial variability across tokens: some tokens concentrate most of their weights on a few experts and are thus robust to expert pruning, whereas others distribute weights more evenly and are therefore more sensitive.
(a) Layer-wise pruning sensitivity.
(b) Token-wise pruning sensitivity.
Figure 5: Sensitivity analysis of MoE experts: (a) layer-wise and (b) token-wise. Both dimensions exhibit large variance, motivating a dynamic pruning strategy.
Dynamic pruning.
Before combining layer‑ and token‑level information, we normalize their sensitivity scores to the range $[0,1]$. Here, $W_{l}$ denotes the KL divergence of layer $l$ when reducing expert selection from $k=8$ to $k=3$. $W_{\min}$ and $W_{\max}$ are defined as the divergences of the least‑sensitive and most‑sensitive layers, respectively. For tokens, $R_{i}$ represents the top‑3/top‑8 routing weight ratio of token $i$ at a specific layer during inference, with $R_{\min}$ and $R_{\max}$ similarly estimated from the calibration set. The normalized sensitivity scores are then defined as:
$L^{\prime}_{l}=(W_{l}-W_{\min})/(W_{\max}-W_{\min}),\qquad T^{\prime}_{i}=(R_{\max}-R_{i})/(R_{\max}-R_{\min}).$
The combined sensitivity score for token $i$ at layer $l$ is then:
$S_{i,l}=\lambda\cdot\tfrac{1}{2}\big(L^{\prime}_{l}+T^{\prime}_{i}\big),\quad\lambda\in(0,1)$
where we simply average layer- and token-level sensitivities to balance their contributions, and use $\lambda$ to control the aggressiveness of pruning. The dynamic number of selected experts is determined as:
$K_{i,l}=\left\lfloor K_{\min}+(K_{\text{base}}-K_{\min})\cdot S_{i,l}\right\rceil.$
Here $K_{\text{base}}=8$ is the default number of experts, and $K_{\min}=3$ serves as a safeguard to prevent selecting too few experts: from empirical inspection, we found that reducing the number of selected experts below this threshold leads to a complete collapse of the model. Details of the experimental validation for the setting of $K_{\min}$ and the choice of $\lambda$ are provided in the Appendix C.
6 Experiments
In this section, we first introduce the experimental setup, including models, datasets, and implementation details. We then separately compare our proposed Pick and Ban approaches with existing methods, highlighting their respective strengths. After that, we demonstrate how combining Ban&Pick achieves both improved performance and faster inference. Finally, we examine the influence of key experts on tasks from other domains, offering insights for future research directions.
6.1 Experiment Setup
Models and Datasets
We conduct experiments on four fine-grained MoE models from two different model families, covering a wide range of parameter scales: DeepSeek-V2-Lite-Chat (16B), DeepSeek-V2.5-1210 (236B), Qwen3-30B-A3B, and Qwen3-235B-A22B. Our evaluation spans three categories of tasks using five commonly adopted datasets: AIME2024 (patel2024aimeaioptimizationmultiple, ) and Math-500 for mathematical reasoning, GPQA-Diamond (rein2023gpqagraduatelevelgoogleproofqa, ) for general knowledge reasoning, and LiveCodeBench (V1) (jain2024livecodebench, ) with HumanEval+ (evalplus, ) for code generation. For DeepSeek-V2-Lite-Chat, we replace AIME2024 with GSM8K, as AIME2024 is overly difficult for this smaller model. We use OpenCompass (2023opencompass, ) as evaluation toolkit. The detailed generation hyperparameters are provided in the Appendix D.
Table 1: Comparison of Pick against three baselines across four models. Reported accuracies cover multiple benchmarks, with GSM8K only for DeepSeek-V2-Lite-Chat.
Models Methods GSM8K/AIME24 Math500 GPQA Livecodebench HumanEval+
Deepseek-V2-Lite-Chat Baseline 66.34 26.40 24.75 9.25 50.61
Dynamic 66.56 26.80 24.75 9.00 50.61
Tips 66.87 26.60 25.25 9.50 51.22
Pick 67.48 28.00 25.76 10.50 51.83
Deepseek-v2.5-1210 Baseline 23.33 81.00 47.47 60.75 83.54
Dynamic 24.00 81.20 46.97 61.00 83.54
Tips 24.67 80.60 47.98 61.25 84.15
Pick 27.34 82.80 49.49 62.75 84.76
Qwen3-30B-A3B Baseline 80.67 95.20 65.66 91.75 71.34
Dynamic 82.66 94.80 66.16 92.00 71.95
Rice 83.33 95.60 67.68 93.50 72.56
Pick 85.33 96.60 69.19 94.25 75.61
Qwen3-235B-A22B Baseline 84.67 96.00 71.21 92.75 78.05
Dynamic 85.33 96.00 71.72 93.00 78.66
Rice 86.67 96.20 72.22 93.00 79.88
Pick 88.00 96.80 74.24 94.50 81.10
Implementation Details
For the Pick method, we follow the procedures described in  Section 4 to identify key experts for each of the three task categories–math, code and general reasoning–and apply targeted enhancements to them. We then evaluate the resulting models on their corresponding task types. For the Ban method, we set $K_{\min}=3$ and $\lambda=0.7$ for all four models, aiming to achieve a balanced trade-off between accuracy and speedup (further discussed in  Appendix C). When combining Pick and Ban, we note that in layers containing key experts, Ban has already pruned redundant experts, so directly applying the range-based replacement strategy (strategy D) leads to noticeable performance degradation. To address this, we instead adopt the range-based addition strategy (strategy C).
6.2 Results of Pick
We compare our proposed Pick method against three representative baselines, covering both routing-strategy adjustments and test-time scaling approaches (see Appendix E for reproducibility details):
(1) Dynamic Routing. (huang-etal-2024-harder, ) adaptively allocates experts according to routing weights, assigning more experts to harder tokens and fewer to easier ones, realized by accumulating experts until a threshold is reached. We set thresholds $\{0.7,0.8,0.9\}$ and report the best result.
(2) Tip. (wang2025thoughtsplaceunderthinkingo1like, ) addresses the issue of underthinking, where models abandon reasoning paths too early. It penalizes tokens that typically indicate switching paths, thereby encouraging the model to continue deepening its current line of reasoning.
(3) Rice. (wang2025expertsneedsteeringthinking, ) leverages explicit <think> tokens to identify cognitive experts that specialize in reasoning. During inference, the routing weights of these experts are multiplied by a constant factor to amplify their influence. Since RICE relies on explicit <think> tokens and is therefore inapplicable to the two DeepSeek models, and it has reported better results than Tip on Qwen3 series, we adopt Tip as the baseline for DeepSeek and RICE as the baseline for Qwen3 in our experiments.
As shown in Table 1, Pick achieves consistent and notable improvements across four models with different parameter scales on five datasets, outperforming all three baselines. The gains are especially pronounced on the Qwen3 series, which we attribute to their stronger expert specialization. Remarkably, on the large-scale Qwen3-235B-A22B, whose performance is already comparable to leading closed-source models, Pick further boosts accuracy by 3.33%, 3.03%, and 3.05% on AIME2024, GPQA, and HumanEval+, respectively. These results highlight the potential of leveraging domain-key experts to further improve MoE model accuracy on diverse tasks. Meanwhile, on DeepSeek-v2.5-1210, Pick still achieves an average improvement of 1.97% across the five datasets, demonstrating strong generalization capability.
6.3 Results of Ban
Table 2: Comparison of Ban against DES and ODP across four models. We report average experts per token (Avg-Topk), accuracy, and inference speedup under the vLLM (0.9.1) backend.
Models Methods Avg-Topk GSM8K/AIME24 Math500 GPQA Livecodebench HumanEval+ Speedup
Deepseek-V2-Lite-Chat Baseline 6.00 66.34 26.40 24.75 9.25 50.61 1.00
DES 4.18 63.76 21.40 22.22 7.25 45.73 1.20
ODP 4.22 64.76 23.20 23.23 7.75 46.95 1.19
Ban 4.02 66.11 25.00 24.24 8.50 48.17 1.22
Deepseek-v2.5-1210 Baseline 6.00 23.33 81.00 47.47 60.75 83.54 1.00
DES 4.21 22.66 80.20 39.90 58.00 79.27 1.21
ODP 4.29 23.33 80.20 43.43 59.25 81.71 1.20
Ban 3.97 23.33 80.60 45.96 60.25 82.93 1.24
Qwen3-30B-A3B Baseline 8.00 80.67 95.20 65.66 91.75 71.34 1.00
DES 4.94 73.33 92.60 56.06 84.50 65.24 1.25
ODP 4.98 74.67 93.20 60.10 86.00 67.07 1.25
Ban 4.82 80.00 94.60 64.65 90.00 70.12 1.25
Qwen3-235B-A22B Baseline 8.00 84.67 96.00 71.21 92.75 78.05 1.00
DES 4.89 68.00 94.80 62.12 90.75 70.12 1.26
ODP 4.91 73.33 95.20 66.16 91.00 73.78 1.26
Ban 4.77 83.33 95.40 69.70 91.25 75.61 1.27
We compare our proposed Ban method against two mainstream dynamic expert-pruning approaches:
(1) DES. (lu-etal-2024-experts, ) (Dynamic Expert Skip) dynamically prunes low-weight experts by observing that, for easy tokens, the contribution of lower-ranked experts is often negligible. Concretely, it decides whether to drop the secondary expert based on the ratio between expert routing weights.
(2) ODP. (huang2024mc, ) (Online Dynamic Pruning) extends DES by adding a key-token protection mechanism. Tokens that receive disproportionately high attention scores are treated as key tokens, and their experts are exempted from pruning to preserve accuracy.
In terms of accuracy, as shown in Table 2, both DES and ODP suffer from severe degradation on these challenging tasks, with drops exceeding 5% on certain datasets. In contrast, Ban limits the accuracy loss to within 1.5% on most models and datasets, except HumanEval+, which we found sensitive to pruning, and even within 1% on more than half of the results. Notably, on Qwen3-30B-A3B, Ban reduces accuracy on AIME2024 by only 0.67%–equivalent to misanswering just one additional problem out of five evaluations–and incurs merely 1.01% degradation on GPQA.
[…]
6.4 Results of Ban&Pick
Table 3: Results of combining Ban&Pick across four models. We report average experts per token, accuracy, and inference speedup under vLLM. Relative changes from baseline are annotated in gray.
Models Methods Avg-Topk GSM8K/AIME24 Math500 GPQA Livecodebench HumanEval+ Speedup
Deepseek-V2-Lite-Chat Baseline 6.00 66.34 26.40 24.75 9.25 50.61 1.00
Ban&Pick 4.03 66.94 ↑0.60 26.80 ↑0.40 25.25 ↑0.50 9.50 ↑0.25 50.00 ↓0.61 1.22
Deepseek-v2.5-1210 Baseline 6.00 23.33 81.00 47.47 60.75 83.54 1.00
Ban&Pick 3.99 26.00 ↑2.67 81.60 ↑0.60 47.98 ↑0.51 61.25 ↑0.50 84.15 ↑0.61 1.23
Qwen3-30B-A3B Baseline 8.00 80.67 95.20 65.66 91.75 71.34 1.00
Ban&Pick 4.83 84.66 ↑3.99 95.80 ↑0.60 68.18 ↑2.52 92.85 ↑1.10 72.56 ↑1.22 1.25
Qwen3-235B-A22B Baseline 8.00 84.67 96.00 71.21 92.75 78.05 1.00
Ban&Pick 4.79 86.67 ↑2.00 96.00 ↑0.00 72.73 ↑1.52 93.25 ↑0.50 78.66 ↑0.61 1.26
[…]
7 Interaction Among Key Experts Across Tasks
We further investigate the scenario where all identified key experts across math, general reasoning, and code are enhanced simultaneously. Using Qwen3-30B-A3B as a case study, we report in Table 4 the performance when enhancing key experts for a single domain, for domain pairs (Math+General, Math+Code, General+Code), and for all three domains together. This setting allows us to examine whether experts from different domains complement or interfere with each other when activated jointly, providing a more comprehensive understanding of cross-domain interactions.
Table 4: Task accuracies (%) on Qwen3-30B-A3B under different combinations of enhanced experts. A ✓ indicates that the corresponding domain experts are enhanced, while × means not enhanced.
Math General Code AIME2024 MATH-500 GPQA-Diamond LiveCodeBench HumanEval+
$\times$ $\times$ $\times$ 80.67 95.20 65.66 91.75 71.34
$\times$ $\times$ 85.33 ↑4.66 96.60 ↑1.40 65.15 ↓0.51 91.50 ↓0.25 71.95 ↑0.61
$\times$ $\times$ 82.00 ↑1.33 95.20 ↑0.00 69.19 ↑3.53 91.75 ↑0.00 72.56 ↑1.22
$\times$ $\times$ 81.33 ↑0.66 95.00 ↓0.20 65.66 ↑0.00 94.25 ↑2.50 75.61 ↑4.27
$\times$ 84.67 ↑4.00 96.20 ↑1.00 67.68 ↑2.02 92.00 ↑0.25 71.34 ↑0.00
$\times$ 84.67 ↑4.00 96.00 ↑0.80 66.16 ↑0.50 93.50 ↑1.75 74.39 ↑3.05
$\times$ 80.67 ↑0.00 94.80 ↓0.40 68.18 ↑2.52 92.75 ↑1.00 75.00 ↑3.66
84.00 ↑3.33 96.00 ↑0.80 67.68 ↑2.02 93.00 ↑1.25 74.39 ↑3.05
From Table 4, we observe that enhancing multiple domains simultaneously still brings clear improvements. For instance, when both Math and General experts are enhanced, AIME2024 increases by +4.0% and GPQA by +2.02%. However, these gains are slightly smaller than enhancing each domain individually ($0.66\%$ and $1.51\%$, respectively). When all three domains are enhanced together, the model continues to outperform the baseline—+3.33% on AIME2024, +2.02% on GPQA, and +3.05% on HumanEval+—though the gains are somewhat reduced compared to the best single-domain setting. These results indicate that while domain-specific key experts do interfere to some extent, their combined enhancement remains clearly beneficial, confirming their crucial role.
8 Conclusion and Implications for Future Work.
In this study, we reveal that in fine-grained MoE models, relatively balanced routing strategy underutilizes a small set of key experts which have a disproportionately large impact on task performance, and fixed expert activation introduces substantial redundancy. By selectively increasing the utilization of key experts and dynamically pruning redundant ones, we release the potential of routing and provide a novel and effective framework that improves both accuracy and efficiency. For future work, we suggest exploring the underlying mechanisms of how key experts enhance model performance, and investigating more sophisticated enhancement strategies beyond uniform boosting.
Appendix A Complete Results on Expert Specialization
In Section 4.1, we analyzed the specialization of experts in fine-grained MoE models using Qwen3-30B-A3B as a case study, and presented representative figures for a subset of layers. Here, we provide the complete visualization of expert usage frequencies across all layers of the model. Results are shown for three representative task types—math, code, and general. For each layer, the bar charts display the relative usage frequency of each expert, aggregated over all tokens in both the prefill and decode stages. Figures are grouped every 12 layers for readability. These results (Figures 69) confirm that different task types consistently activate distinct subsets of experts, highlighting clear patterns of domain specialization.
Figure 6: Expert usage frequency for layers 0–11 across math, code, and general tasks.
Figure 7: Expert usage frequency for layers 12–23 across math, code, and general tasks.
Figure 8: Expert usage frequency for layers 24–35 across math, code, and general tasks.
Figure 9: Expert usage frequency for layers 36–47 across math, code, and general tasks.
Appendix B Implementation details of Top-1000 KL divergence
Following prior work [49], we note that the model’s output logits typically follow a long-tail distribution, where only the top-ranked tokens carry substantial probability mass. Therefore, when computing KL divergence, we restrict attention to the Top-1000 tokens. Concretely, let $p$ denote the probability distribution obtained from the original model, and $q$ the distribution when a candidate expert is pruned. Denote by $\mathcal{I}_{1000}$ the index set of the top-1000 tokens under $p$. The restricted distributions are defined as
$p^{\prime}(i)=\frac{p(i)}{\sum_{j\in\mathcal{I}_{1000}}p(j)},\quad q^{\prime}(i)=\frac{q(i)}{\sum_{j\in\mathcal{I}_{1000}}q(j)},\quad i\in\mathcal{I}_{1000}.$
The KL divergence used in our analysis is then
$D_{\mathrm{KL}}(p^{\prime}\parallel q^{\prime})=\sum_{i\in\mathcal{I}_{1000}}p^{\prime}(i)\log\frac{p^{\prime}(i)}{q^{\prime}(i)}.$
This definition ensures that the comparison focuses on the most influential tokens, while ignoring the long tail of near-zero probabilities.
Appendix C Hyperparameter Choices in Dynamic Pruning
As introduced in Section 5, our dynamic pruning strategy involves three hyperparameters: $\beta$, $K_{\min}$, and $\lambda$. The parameter $\beta$ balances the contributions of layer-level and token-level sensitivities. Since our motivation is to jointly consider both factors in a balanced way, we simply set $\beta=0.5$ empirically and do not explore it further. In this appendix, we therefore focus on discussing the selection of $K_{\min}$ and $\lambda$ in detail.
Choice of $K_{\min}$.
In the main text, we noted that reducing the number of selected experts below a certain threshold causes the model to collapse entirely. Therefore, $K_{\min}$ must guarantee that even when both layer- and token-level sensitivities are low, the number of selected experts never falls below this safe lower bound.
To study this effect, we take Qwen3-30B-A3B as an example and directly enforce the number of experts per token to be fixed between 2 and 7 (with $K_{\text{base}}=8$ as the default). Table 5 reports the performance on five benchmarks. For efficiency, results on AIME2024 (except the baseline) are measured over a single run.
TopK AIME2024 Math500 GPQA Livecodebench HumanEval+
8 80.67 95.2 65.66 91.75 71.34
7 80.00 95.0 65.66 91.25 71.95
6 76.67 93.2 63.64 89.75 69.51
5 76.67 88.2 59.60 87.00 66.46
4 63.33 78.4 54.55 74.75 59.76
3 40.00 55.8 44.95 57.50 43.29
2 16.67 19.6 21.72 25.50 17.07
Table 5: Performance of Qwen3-30B-A3B under different fixed numbers of selected experts.
We observe that removing just one expert (TopK=7) leads to almost no performance degradation, consistent with the expert pruning observations reported in Section 3 in the main text. However, once two or more experts are pruned, noticeable drops appear across all benchmarks. The most critical boundary occurs at TopK=3: performance collapses when fewer than three experts are selected. Therefore, we set $K_{\min}=3$ to prevent the model from falling below this threshold. On the DeepSeek series, we observed a similar phenomenon, further reinforcing this choice.
Choice of $\lambda$.
The hyperparameter $\lambda$ controls the trade-off between accuracy preservation and inference speedup: larger values of $\lambda$ bias the pruning strategy toward retaining more experts (thus preserving accuracy), while smaller values encourage more aggressive pruning (thus improving throughput). To study this trade-off, we take Qwen3-30B-A3B as an example and vary $\lambda$ from $0.5$ to $0.9$ in increments of $0.1$. For each setting, we evaluate accuracy on the same five benchmarks as in the main experiments and measure the inference speedup using vLLM as the serving backend. The speedup is calculated as the ratio between the total inference time across all datasets and the baseline time without pruning. All results are reported from a single run.
$\lambda$ Avg TopK AIME2024 Math500 GPQA Livecodebench HumanEval+ Speedup
8 80.67 95.2 65.66 91.75 71.34 1.00
0.5 4.33 73.33 93.8 59.60 86.50 65.85 1.30
0.6 4.61 76.67 94.0 63.64 89.25 68.90 1.26
0.7 4.82 80.00 94.6 64.65 90.00 70.12 1.25
0.8 5.09 80.00 94.0 64.65 90.75 70.12 1.22
0.9 5.32 80.00 94.8 64.14 90.50 70.73 1.20
Table 6: Accuracy and inference speedup of Qwen3-30B-A3B under different values of $\lambda$. The first row corresponds to the baseline.
From Table 6, we observe that when $\lambda$ is below $0.7$, the model achieves higher speedups but suffers noticeable accuracy degradation compared with $\lambda=0.7$. When $\lambda$ exceeds $0.7$, accuracy improves slightly, but the gains are marginal relative to the loss in efficiency. Considering this trade-off, we set $\lambda=0.7$ in our main experiments and report results under this configuration. Nevertheless, $\lambda$ remains a tunable knob that allows users to flexibly balance accuracy preservation and inference acceleration according to their own application needs.
Appendix D Evaluation Details
D.1 Dataset Evaluation.
For AIME2024, due to the small sample size, we follow the common practice in technical reports by evaluating each method five times and reporting the average. For other datasets, constrained by time, we conduct only a single evaluation. On the GPQA-Diamond dataset, to further reduce errors caused by misaligned answer formats, we adopt an LLM-judge in addition to the base evaluator: specifically, we deploy a Qwen2.5-72B-Instruct model [33] with vLLM [34] as the judge model. For all other datasets, we rely on the default base evaluators provided by OpenCompass to compute accuracy.
D.2 Generation Hyperparameters.
During generation, for all four models used in our experiments, we adopt the recommended hyperparameters provided in their original papers or Model Cards. The configuration is summarized in Table 7.
Table 7: Generation hyperparameters for the evaluated models.
Model Top-k Top-p Temperature Max Output Length
DeepSeek-V2-Lite-Chat None 0.95 0.3 8192
DeepSeek-V2.5-1210 1 0.90 0.3 8192
Qwen3-30B-A3B 20 0.95 0.6 32768
Qwen3-235B-A22B 20 0.95 0.6 32768
Appendix E Details of Compared Methods and Reproduction Settings
In the main experiments, we compared our proposed Ban and Pick strategies against several existing baselines. For brevity, the main text only provided a concise overview of these methods. In this appendix, we present a more detailed description of each compared method, together with the specific reproduction settings used in our experiments. We organize the discussion into two parts: methods compared with Pick, and methods compared with Ban.
E.1 Compared Methods for Pick
Dynamic Routing.
[18] first proposed the idea of dynamically allocating experts, arguing that assigning a fixed number of experts to each token is overly rigid. Instead, the number of experts should be adapted according to the difficulty of the underlying task: tokens from more challenging tasks should be allocated more experts, while tokens from easier tasks can be served with fewer. Concretely, experts are sequentially selected in descending order of routing weight until the cumulative weight exceeds a predefined threshold $\tau$. Formally, let $r_{(1)},r_{(2)},\dots,r_{(K)}$ denote the routing weights of experts sorted in descending order. The number of selected experts $k$ is determined as
$k=\min\Bigl\{m\,\big|\,\sum_{j=1}^{m}r_{(j)}\geq\tau\Bigr\}.$
In our experiments, since the optimal threshold $\tau$ cannot be known in advance, we evaluate three candidate values $\tau\in\{0.7,0.8,0.9\}$ and report in the main text the result corresponding to the setting that achieves the best average accuracy across benchmarks.
TiP.
The core idea of TiP is motivated by the observation that, during chain-of-thought (CoT) reasoning, models often switch reasoning paths too frequently in the early exploration stage. Such premature abandonment of a correct path is referred to as underthinking. To mitigate this issue, TiP introduces a penalty on tokens that typically indicate a switch of reasoning path (e.g., “alternatively”). Concretely, after the softmax operation, the probability mass of these tokens is reduced by subtracting a fixed penalty value from their logits, thereby encouraging the model to continue deepening its current reasoning path rather than switching too early.
This method involves two hyperparameters: $\alpha$ (penalty strength) and $\beta$ (penalty duration). The original paper did not provide an automatic tuning strategy but instead conducted a grid search to identify relatively effective values. Following their reported settings, we set $\alpha=3$ and $\beta=600$ in our reproduction.
RICE.
RICE builds on the observation that recent MoE-LLMs often employ an explicit <think> token to train chain-of-thought (CoT) reasoning ability. For example, the Qwen3 series can switch between deep-thinking and non-thinking modes by toggling the use of the <think> token. RICE leverages this property by comparing expert activation frequencies with and without the <think> token, thereby identifying cognitive experts that are closely associated with deep reasoning. Once identified, the routing weight of a cognitive expert is multiplied by a constant factor greater than $1$ whenever it is selected, amplifying its influence and enhancing the model’s reasoning ability.
Since this method relies on explicit <think> tokens, it is only applicable to models such as the Qwen3 series, but not to other models used in our experiments. For reproduction, we followed the original paper’s settings: for Qwen3-235B-A22B, we directly used the cognitive experts reported therein; for Qwen3-30B-A3B, as the original paper did not list them, we applied the ICE (Identify Cognitive Experts) procedure proposed in the paper to locate the corresponding cognitive experts before applying the enhancement.
E.2 Compared Methods for Ban
DES.
[…]
In our reproduction, we aimed to align the pruning degree and acceleration ratio with our proposed Ban method, so that the comparison mainly reflects differences in accuracy preservation. For Qwen3 models, when $\lambda=0.7$, the average number of selected experts under Ban is slightly below 5. Therefore, we set the maximum number of experts to 6 and the minimum to 4. Specifically, we computed the medians of the ratios $\tfrac{r_{(4)}}{r_{(5)}}$ and $\tfrac{r_{(5)}}{r_{(6)}}$ on a calibration set, where $r_{(j)}$ denotes the $j$-th largest routing weight. At inference time, if the first ratio exceeds its median, only the top-4 experts are kept; otherwise, if the first ratio does not exceed its median but the second does, the top-5 experts are kept; otherwise, all top-6 experts are retained. For the DeepSeek models, we followed an analogous procedure to ensure comparable pruning degree and speedup.
ODP.
ODP(Online Dynamic Pruning) is an extension of DES that incorporates an additional key-token protection mechanism. While DES prunes experts dynamically based on weight ratios, ODP further observes that certain tokens carry disproportionately high importance in the attention map. If a token receives an attention score significantly larger than the others, pruning its experts may harm model performance. To address this, ODP protects such key tokens: even if their expert weights satisfy the pruning condition, they are exempted from pruning.
In our reproduction, we followed exactly the same procedure as DES for the pruning part. On top of that, we adopted the evaluation metric for token importance proposed in the ODP paper to identify key tokens. Whenever a token is classified as key, we enforced the selection of all 6 experts for it, regardless of the pruning condition.